OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

i wil help u if u help me

OpenStudy (anonymous):

Use estimation to determine which of the following quotients you expect to be less than 1. four-fifteenths divided by six-fifteenths seven-eighths divided by three-eighths nine-elevenths divided by one-eleventh five-sevenths divided by one-seventh

OpenStudy (nurali):

|x + 2| + 16 = 14 |x + 2| =-2 x+2=-2 or x+2=2 x=-4 or x=0

OpenStudy (anonymous):

thats not the right answer Nurali

OpenStudy (nurali):

Second option is correct.

OpenStudy (anonymous):

no its not, i put it on my test and it said it was wrong..

OpenStudy (jdoe0001):

hmm is it |x + 2| - 16 = 14 ? maybe for the one you posted, Nurali's line are correct

OpenStudy (anonymous):

i clicked that on my test and i got 0 points for it, which means its wrong.

OpenStudy (jdoe0001):

anyhow, maybe you have a typo or something

OpenStudy (anonymous):

i copied and pasted.

OpenStudy (jdoe0001):

well, the answer provided by Nurali is correct

OpenStudy (anonymous):

WELL NOT ACCORDING TO MY TEST

OpenStudy (jdoe0001):

then the test is wrong =)

OpenStudy (anonymous):

OK!

thomaster (thomaster):

Also, you are cheating on a test, which is strictly prohibited here on OpenStudy.

OpenStudy (anonymous):

oh shutup